11.An ultrasound transducer is used to direct longitudinal waves at a frequency of 12MHz at an artery . The waves reflecting back to the transducer at an angle of 340 to the direction of blood flow. The blood cells are moving at a speed of 0.31ms-1 in the artery, and the speed of sound in the area is 1550ms-1. Calculate the change in frequency of the wave you would expect to see due to the Doppler effect. 


1
Expert's answer
2021-03-29T09:04:27-0400
"f_e=\frac{c+v}{c}f_0\\
f_e=\frac{1550+0.31}{1550}12000000=12002400\ Hz\\\Delta f=2400\ Hz"